OpenStudy (anonymous):

Given the system of equations, what is the solution? x + 3y = 5 x - 3y = -1 {(-4, -3)} {(2, 1)} {(-2, 3)}

OpenStudy (anonymous):

y= x/3 + 1/3 x + 3(x/3 + 1/3)=5 2x+1=5 x=2 then just solve for y